# Demons registration for 2D empirical wavelet transform: Application to texture segmentation

### v2

@@ Abstract
The empirical wavelet transform is a fully adaptive time-scale representation that has been widely used in the last decade. Inspired by the empirical mode decomposition, it consists of filter banks based on harmonic mode supports. Recently, it has been generalized to build the filter banks from any generating function using mappings. In practice, the harmonic mode supports can have low constrained shape in 2D, leading to numerical difficulties to compute the mappings and therefore the related wavelet filters. This work aims to propose an efficient numerical scheme to compute empirical wavelet coefficients using the demons registration algorithm. Results show that the proposed approach gives a numerically robust wavelet transform. An application to texture segmentation of scanning tunnelling microscope images is also presented.
@@ 1 Introduction
The empirical wavelet transform is a fully adaptive time-scale representation, introduced in [5], based on data-driven filter banks. Unlike the traditional wavelet transform, the wavelet filters are not designed on the basis of prescribed scales independent of the data under study, but on the basis of supports that contain the underlying harmonic modes. This approach is inspired by the empirical mode decomposition [13], but is based on solid theoretical foundations that the latter lacks. Due to its robust performance in decomposing images, it has led to many applications such as glaucoma detection [19, 21], hyperspectral image classification [20], cancer histopathological image classification [4], medical image fusion [23, 22] and texture segmentation [15]. Notably, it has been shown to outperform traditional wavelet transforms in extracting appropriate texture features from images [14, 16].
@@
The last decade has seen an intensive development of the 2D empirical wavelet transform. The extraction of harmonic modes can be performed by scale-space representations [7] and several methods of separation of the harmonic modes in possibly low constrained shaped supports have been proposed, such as the Tensor [5], Ridgelet, Curvelet [16], Watershed [9] and Voronoi [6] methods. However, the design of filter banks has mainly long been limited to a specific generating function based on the Littlewood-Paley formulation. This formulation relies on a separable definition or the distance to the boundaries of the supports, which limits its extension to other generating functions. Recently, a general framework has been proposed to consider classic generating functions [18]. Deformations of a homogeneous or separable wavelet kernel are carried out by mappings between the data’s Fourier supports and the generating function’s Fourier support. In practice, this approach suffers from the difficulty to estimate the required mappings with constraints of invertibility, continuity and differentiability.
In this work, we propose a numerical implementation of the empirical wavelet transform based on the demons algorithm, widely used in image registration. We particularly compare different variants of the demons algorithm for the design of wavelet filters and the wavelet reconstruction from homogeneous and separable wavelet kernels. An application to the texture segmentation of scanning tunnelling microscope images is also considered for illustration. In Section 3, we recall the framework of empirical wavelet transform based on mappings, introduce specific band-pass wavelet filters and present different demons algorithms for mapping estimation. In Section 4, we compare numerically the mapping estimators for empirical wavelet transforms and show the application to texture segmentation. Finally, a discussion is provided in Section 5. A Matlab toolbox will be made publicly available at the time of publication.
@@ 3.1 Empirical wavelet systems
The coonstruction of empirical wavelets relies on the partitioning of the Fourier domain $\Omega$. Technically, we consider a family of disjoint open sets $\{\Omega_{n}\}_{n\in\Upsilon}$, with $\Upsilon\subset{\mathbb{Z}}$, of closures $\overline{\Omega}_{n}$ covering the Fourier domain, i.e., $\Omega=\bigcup_{n\in\Upsilon}\overline{\Omega_{n}}$. In this work, we focus on real-valued images, which implies that the Fourier domain is symmetric, leading to consider a partition $\{\Omega_{n}\}_{n\in\Upsilon}$ that is symmetric, i.e., such that $\Omega_{-n}=\{-\xi\mid\xi\in\Omega_{n}\}$. To obtain such a partition, the modes of the Fourier spectrum are obtained by scale-space representation [7] and the Fourier domain is partitioned by the Watershed [16] or Voronoi [6] methods. Figure 1 shows an example of an image and the Watershed and Voronoi partitions of the logarithm of its Fourier spectrum using a scale-space step-size parameter set to $s_{0}=0.8$.
@@
in most cases (see [18] for more details). If $\gamma_{n}$ is not a diffeomorphism, the substitution theorem ensuring the energy conservation (1) is not valid despite the use of the normalization coefficient $a_{n}(\xi)$. However, in practice, estimating diffeomorphisms is a difficult and computationally expensive task. Since an intervertible continuous function is a homeopmorphism if and only if it is an open map, i.e., a map for which the preimage of an open set is an open set, the existence of homeomorphisms is a milder assumption and their estimation is less difficult. We thus propose an unnormalized definition for the case when $\gamma_{n}$ is only a homeomorphism by removing the Jacobian determinant from Definition 1 as follows.
@@ 3.2 Band-pass empirical wavelets
For the study of the numerical behavior of the empirical wavelet transform, it is more convenient to consider band-pass filters. Therefore, we propose in this section to define homogeneous or separable 2D wavelet band-pass filters. The band-pass wavelet filters proposed in [16, 5] relie on the semi-Euclidean distance to the boundaries of the supports, which is equivalent to a discrete mapping. In 1D, this wavelet kernel can be adapted to the continuous framework as follows.
@@
The square band-pass wavelet, mostly supported by the square $\Lambda=[-\frac{1}{2},\frac{1}{2}]\times[-\frac{1}{2},\frac{1}{2}]$, is defined by, for every $\xi=(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}$,
@@
Let $\chi^{\mathrm{D},\tau}_{n}=\chi^{\mathrm{D}}_{\tau}\circ\gamma_{n}$ and $\chi^{\mathrm{S},\tau}_{n}=\chi^{\mathrm{S}}_{\tau}\circ\gamma_{n}$ the (normalized or unnormalized) symmetric wavelet systems resulting from the disk and square band-pass wavelets $\widehat{\psi}_{\tau}^{D}$ and $\widehat{\psi}_{\tau}^{S}$, respectively, for every $n\in\Upsilon$. The reconstruction formula (2) holds for the normalized and unnormalized systems $\{\chi^{\mathrm{D},\tau}_{n}\}_{n\in\Upsilon}$ and $\{\chi^{\mathrm{S},\tau}_{n}\}_{n\in\Upsilon}$.
@@
Let $\widehat{\psi}_{\tau}$ be either $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{D}$ or $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{S}$. We set $\Lambda_{\tau}=\mathrm{B}(0,\frac{1}{2}+\tau)$ if $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{D}$ and $\Lambda_{\tau}=[-\frac{1}{2}-\tau,\frac{1}{2}+\tau]\times[-\frac{1}{2}-\tau,\frac{1}{2}+\tau]$ if $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{S}$. Let $\xi\in\Omega_{m}$. Since there exists a finite number of $\gamma_{n}$ such that $\gamma_{n}(\xi)\in\Lambda_{\tau}$, we have
$\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{n}(\xi)\right|^{2}=\sum_{n\in\Upsilon^{+}}\left|\frac{1}{a_{n}(\xi)}\right|\left|\widehat{\psi}_{\tau}(\gamma_{n}(\xi))\right|^{2}=\sum_{n\in\Upsilon_{m}^{+}}\left|\frac{1}{a_{n}(\xi)}\right|\left|\widehat{\psi}_{\tau}(\gamma_{n}(\xi))\right|^{2}<+\infty,$
where $\Upsilon_{m}^{+}=\{n\in\Upsilon^{+}:\gamma_{n}(\xi)\in\Lambda_{\tau}\}$ and $a_{n}(\xi)=1$ or $a_{n}(\xi)=1/|\mathrm{det}\;J_{\gamma_{n}}(\xi)|$ for the normalized or unnormalized empirical walet system, respectively. Moreover, since $\widehat{\psi}_{\tau}(\gamma_{m}(\xi))\neq 0$ and $a_{n}(\xi)>0$ for every $n\in\Upsilon$, we have $0<\sum_{n\in\Upsilon^{+}}|\widehat{\psi}_{n}(\xi)|^{2}$. Similarly, we can show that $\sum_{n\in\Upsilon^{+}}|\widehat{\psi}_{n}(\xi)\widehat{\psi}_{-n}(\xi)|<+\infty$. Hence the reconstruction formula (2) for $\{\chi^{\mathrm{D},\tau}_{n}\}_{n\in\Upsilon}$ and $\{\chi^{\mathrm{S},\tau}_{n}\}_{n\in\Upsilon}$ holds by Theorem 1. ∎
@@ 3.3 Mapping estimation
In practice, the mappings $\gamma_{n}$ in Definition LABEL:def:norm_system,def:unnorm_system have to be estimated. The Thirion’s demons algorithm [24] is a mapping estimation scheme inspired by a diffusion process in which the targeted mapping is represented by a displacement field $\gamma$. This method alternates between solving the flow equations and regularization. To give a theoretical justification to this method, [25] revisited it as the following optimisation problem:
@@
where $\Lambda$ and $\Omega$ are considered as functions from ${\mathbb{R}}^{2}$ to ${\mathbb{R}}$, ${\mathcal{F}}({\mathbb{R}}^{2})$ denotes the set of mappings on ${\mathbb{R}}^{2}$, $c$ is an intermediate mapping to account for noisy data, $\sigma_{x}$ controls the spatial uncertainty on the mappings and $\mathcal{R}(\gamma)$ is a regularization term.
The additive demons proposed in [3] consists of estimating the minimum $s$ of Equation 3 using an intermediate displacement field $u$ such that $c=\gamma+u$ and alternate
@@
where $\nabla$ denotes the gradient of functions. This implies that $\sigma_{x}$ controls the maximum step length: $\|u(p)\|_{2}\leq\sigma_{x}/2$ for every $p\in{\mathbb{R}}^{2}$. As for the minimization of Equation 5 with a regularization term $\mathcal{R}(\gamma)=\|\nabla\gamma\|_{2}^{2}/\sigma_{d}^{2}$, it corresponds to a convolution with a Gaussian kernel $G_{\rm diff}(\sigma_{d})$ of standard deviation $\sigma_{d}$. This regularization $\mathcal{R}(\gamma)$ has been modified by [3] in order to add a fluid-like regularization $G_{\rm fluid}(\sigma_{f})$ of standard deviation $\sigma_{f}$. To ensure the mapping $\gamma$ to be diffeomorphic, an alternative update of $c$ with an exponential field has been proposed in [25]: $c=s\circ\exp(u)$. The additive and diffeomorphic demons are summarized in Algotirhm 1.
@@ 9       Diffusion-like regularization: $\gamma^{[k]}=G_{\rm diff}(\sigma_{d})\star c^{[k]}$ ;
10Output: Displacement field $\gamma^{[K]}$.
Algorithm 1 Vercauteren’s demons
@@ 6       $\gamma^{[k]}$leftarrow Upsample $2^{k}\gamma^{*}$ by $2^{k}$
Algorithm 2 Multiresolution demons
@@ 4 Numerical experiments
In this section, we compare the Thirion’s demons (Matlab function imregdemons) and the additive or diffeomorphic Vercauteren’s demons (Matlab toolbox available at https://www.mathworks.com/matlabcentral/fileexchange/39194-diffeomorphic-log-demons-image-registration) for the computation of estimates $\widetilde{\gamma}_{n}$ of the mappings $\gamma_{n}$. We consider the toy image and the sets $\Omega_{n}$ obtained by either Watershed or Voronoi partitioning presented in Figure 1.
@@ 4.1 Mapping estimation set-up
For the three demons, the diffusion-like parameter $\sigma_{d}$ and number of resolution levels $N_{\rm level}$ are selected by minimizing the quadratic risk $\|\Lambda-\Omega_{n}\circ\widetilde{\gamma}_{n}\|_{2}^{2}$ on the values $(\sigma_{d},N_{level})\in(0.3,0.31,\ldots,0.5)\times(n_{P}-1,n_{P})$, with $n_{P}$ the highest integer such that $2^{n_{P}}$ is smaller than each dimension of the image. For the Thirions’s demons, the numbers of iterations at the $N_{level}$ pyramid levels are set to $(2^{4},\ldots,2^{N_{level}+1})$ from the highest to the lowest pyramid level. For the Vercauteren’s demons, the maximum step length is set to $\sigma_{x}=5$, the fluid-like regularization is set to $\sigma_{f}=1$, the error threshold in the stop criterion is set to $\epsilon=10^{-3}$ and the maximum number of iterations is set to $K=500$.

### v3

# Demons registration for 2D empirical wavelet transforms

@@ Abstract
The empirical wavelet transform is a fully adaptive time-scale representation that has been widely used in the last decade. Inspired by the empirical mode decomposition, it consists of filter banks based on harmonic mode supports. Recently, it has been generalized to build the filter banks from any generating function using mappings. In practice, the harmonic mode supports can have a low-constrained shape in 2D, leading to numerical difficulties to estimate mappings adapted to the construction of empirical wavelet filters. This work aims to propose an efficient numerical scheme to compute empirical wavelet coefficients using the demons registration algorithm. Results show that the proposed approach is robust, accurate, and continuous wavelet filters permitting reconstruction with a low signal-to-noise ratio. An application for texture segmentation of scanning tunneling microscope images is also presented.
@@ 1 Introduction
The empirical wavelet transform is a fully adaptive time-scale representation, introduced in [6], based on data-driven filter banks. Unlike the traditional wavelet transform, the wavelet filters are not designed on the basis of prescribed scales independent of the data under study, but on the basis of supports that contain the underlying harmonic modes. This approach is inspired by the empirical mode decomposition [14] but is based on solid theoretical foundations that the latter lacks. Due to its robust performance in decomposing images, it has led to many applications such as glaucoma detection [20, 22], hyperspectral image classification [21], cancer histopathological image classification [4], medical image fusion [24, 23] and texture segmentation [16]. Notably, it has been shown to outperform traditional wavelet transforms in extracting appropriate texture features from images [15, 17].
@@
The last decade has seen an intensive development of the two steps involved in the 2D empirical wavelet transform: $(i)$ the extraction of regions containing the harmonic modes and $(ii)$ the design of filters mainly supported by these regions. The extraction of harmonic modes can be performed by scale-space representations [8], and several methods of separation of the harmonic modes in possibly low-constrained shaped supports have been proposed, such as the Tensor [6], Ridgelet, Curvelet [17], Watershed [10] and Voronoi [7] methods. However, the design of filter banks has mainly long been limited to a specific generating function based on the Littlewood–Paley formulation. This formulation relies on a separable definition or the distance to the boundaries of the supports, which limits its extension to other generating functions.
Recently, a general framework has been proposed to consider classic generating functions [19]. Deformations of a homogeneous or separable wavelet kernel are carried out by mappings between the data’s Fourier supports and the generating function’s Fourier support. In practice, this approach suffers from the difficulty of estimating the required mappings with constraints of invertibility, continuity, and differentiability. Such an estimation can be performed by the well-known demons algorithm [26], as first proposed by [19], or its variants [3, 27]. These fast algorithms have been widely used and compared for medical image registration [28, 25], where they have shown to provide robust and accurate estimates. However, an in-depth investigation of their use for empirical wavelet transform remains a key challenge. This is the main objective addressed in this work.
This manuscript is organized as follows. In Section 3, we recall the framework of empirical wavelet transform based on mappings and present different demons algorithms for mapping estimation. In particular, as a first contribution, we build specific band-pass wavelet filters from homogeneous or separable wavelet kernels, which is suitable for an empirical analysis of the wavelet transform. Section 4 presents the main contribution of this manuscript: a comparative analysis of various demons algorithms for designing continuous wavelet filters. The accuracy of the estimated filters and the effect of the estimation process on wavelet reconstruction is thoroughly analyzed. Additionally, to demonstrate the practical utility of the proposed tools, we include an illustration of texture segmentation for scanning tunneling microscope images. Section 4 provides a detailed discussion of the findings. The MATLAB toolbox for the empirical wavelet transform is freely available at https://github.com/jegilles/Empirical-Wavelets.
@@ 3.1 Empirical wavelet systems
The construction of empirical wavelets relies on the partitioning of the Fourier domain $\Omega$. Technically, we consider a family of disjoint open sets $\{\Omega_{n}\}_{n\in\Upsilon}$, with $\Upsilon\subset{\mathbb{Z}}$, of closures $\overline{\Omega}_{n}$ covering the Fourier domain, i.e., $\Omega=\bigcup_{n\in\Upsilon}\overline{\Omega_{n}}$. In this work, we focus on real-valued images, which implies that the Fourier domain is symmetric, leading us to consider a partition $\{\Omega_{n}\}_{n\in\Upsilon}$ that is symmetric, i.e., such that $\Omega_{-n}=\{-\xi\mid\xi\in\Omega_{n}\}$. To obtain such a partition, the modes of the Fourier spectrum are obtained by scale-space representation [8] , and the Fourier domain is partitioned by the Watershed [17] or Voronoi [7] methods. Figure 1 shows an example of an image and the Watershed and Voronoi partitions of the logarithm of its Fourier spectrum using a scale-space step-size parameter set to $s_{0}=0.8$.
Symmetric empirical wavelet systems are built as filters mostly supported on the sets $\Omega_{n}\cup\Omega_{-n}$ for $n\in\Upsilon^{+}=\{n\in\Upsilon\mid n\geq 0\}$. The authors of [19]have proposed building filters $\chi_{n}$ from a wavelet kernel $\psi$ using mappings $\gamma_{n}$ on ${\mathbb{R}}^{2}$. Let $\psi$ be such that its Fourier transform $\widehat{\psi}$ is localized in frequency and compactly supported (or rapidly decaying) by a connected open subset $\Lambda$. Let $\{\gamma_{n}\}_{n\in\Upsilon}$ be mappings on ${\mathbb{R}}^{2}$ such that $\Lambda=\gamma_{n}(\Omega_{n})\textrm{ if }\Omega_{n}$ is bounded and $\Lambda\subsetneq\gamma_{n}(\Omega_{n})$ otherwise. Empirical wavelet systems $\psi_{n}$ and their symmetric counterparts $\chi_{n}$ are then defined as follows.
@@
in most cases (see [19] for more details). If $\gamma_{n}$ is not a diffeomorphism, the substitution theorem ensuring the energy conservation (1) is not valid despite the use of the normalization coefficient $a_{n}(\xi)$. However, in practice, estimating diffeomorphisms is a difficult and computationally expensive task. Since an intervertible continuous function is a homeopmorphism if and only if it is an open map, i.e., a map for which the preimage of an open set is an open set, the existence of homeomorphisms is a milder assumption and their estimation is less difficult. We thus propose an unnormalized definition for the case when $\gamma_{n}$ is only a homeomorphism by removing the Jacobian determinant from Definition 1, as follows.
@@ 3.3 Band-pass empirical wavelets
For the study of the numerical behavior of the empirical wavelet transform, it is more convenient to consider band-pass filters. Therefore, we propose in this section defining homogeneous or separable 2D wavelet band-pass filters. The proposed definition extends the band-pass wavelet filters introduced in [17, 6], which are based on the semi-Euclidean distance to support boundaries, thus corresponding to a discrete mapping. In 1D, the continuous band-pass wavelet kernel can be defined as follows.
@@
The square band-pass wavelet, mostly supported by the square $\Lambda=[-\frac{1}{2},\frac{1}{2}]^{2}$, is defined by, for every $\xi=(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}$,
@@
Let $\chi^{\mathrm{D},\tau}_{n}=\chi^{\mathrm{D}}_{\tau}\circ\gamma_{n}$ and $\chi^{\mathrm{S},\tau}_{n}=\chi^{\mathrm{S}}_{\tau}\circ\gamma_{n}$, the (normalized or unnormalized) symmetric wavelet systems resulting from the disk and square band-pass wavelets $\widehat{\psi}_{\tau}^{D}$ and $\widehat{\psi}_{\tau}^{S}$, respectively, for every $n\in\Upsilon$. The reconstruction Formula (2) holds for the normalized and unnormalized systems $\{\chi^{\mathrm{D},\tau}_{n}\}_{n\in\Upsilon}$ and $\{\chi^{\mathrm{S},\tau}_{n}\}_{n\in\Upsilon}$.
@@
Let $\widehat{\psi}_{\tau}$ be either $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{D}$ or $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{S}$. We set $\Lambda_{\tau}=\mathrm{B}(0,\frac{1}{2}+\tau)$ if $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{D}$ and $\Lambda_{\tau}=[-\frac{1}{2}-\tau,\frac{1}{2}+\tau]^{2}$ if $\widehat{\psi}_{\tau}=\widehat{\psi}_{\tau}^{S}$. Let $\xi\in\Omega_{m}$. Since there exists a finite number of $\gamma_{n}$ such that $\gamma_{n}(\xi)\in\Lambda_{\tau}$, we have
$\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{n}(\xi)\right|^{2}=\sum_{n\in\Upsilon^{+}}\left|\frac{1}{a_{n}(\xi)}\right|\left|\widehat{\psi}_{\tau}(\gamma_{n}(\xi))\right|^{2}=\sum_{n\in\Upsilon_{m}^{+}}\left|\frac{1}{a_{n}(\xi)}\right|\left|\widehat{\psi}_{\tau}(\gamma_{n}(\xi))\right|^{2}<+\infty,$ (3)
where $\Upsilon_{m}^{+}=\{n\in\Upsilon^{+}:\gamma_{n}(\xi)\in\Lambda_{\tau}\}$ and $a_{n}(\xi)=1$ or $a_{n}(\xi)=1/|\mathrm{det}\;J_{\gamma_{n}}(\xi)|$ for the normalized or unnormalized empirical wavelet system, respectively. Moreover, since $\widehat{\psi}_{\tau}(\gamma_{m}(\xi))\neq 0$ and $a_{n}(\xi)>0$ for every $n\in\Upsilon$, we have $0<\sum_{n\in\Upsilon^{+}}|\widehat{\psi}_{n}(\xi)|^{2}$. Similarly, we can show that $\sum_{n\in\Upsilon^{+}}|\widehat{\psi}_{n}(\xi)\widehat{\psi}_{-n}(\xi)|<+\infty$. Hence, the reconstruction Formula (2) for $\{\chi^{\mathrm{D},\tau}_{n}\}_{n\in\Upsilon}$ and $\{\chi^{\mathrm{S},\tau}_{n}\}_{n\in\Upsilon}$ holds by Theorem 1. ∎
@@ 3.4 Mapping estimation
In practice, the mappings $\gamma_{n}$ in Definitions 1 and 2 have to be estimated. The Thirion’s demons algorithm [26] is a mapping estimation scheme inspired by a diffusion process in which the targeted mapping is represented by a displacement field $\gamma$. This method alternates between solving the flow equations and regularization. To give a theoretical justification for this method, [27] revisited it as the following optimization problem:
@@
where $\Lambda$ and $\Omega$ are considered as functions from ${\mathbb{R}}^{2}$ to ${\mathbb{R}}$, ${\mathcal{F}}({\mathbb{R}}^{2})$ denotes the set of mappings on ${\mathbb{R}}^{2}$, $c$ is an intermediate mapping to allow spatial uncertainties, $\sigma_{x}$ is a parameter controlling the spatial uncertainty on the mappings, and $\mathcal{R}(\gamma)$ is a regularization term.
The additive demons algorithm proposed in [3] consists of estimating the minimum $s$ of Equation 4 using an intermediate displacement field $u$ such that $c=\gamma+u$ and alternate
@@
where $\nabla$ denotes the gradient of functions. This implies that $\sigma_{x}$ controls the maximum step length: $\|u(p)\|_{2}\leq\sigma_{x}/2$ for every $p\in{\mathbb{R}}^{2}$. As for the minimization of Equation 6 with a regularization term $\mathcal{R}(\gamma)=\|J_{\gamma}\|_{2}^{2}/\sigma_{d}^{2}$, it corresponds to a convolution with a Gaussian kernel $G_{\rm diff}(\sigma_{d})$ of standard deviation $\sigma_{d}$. This regularization $\mathcal{R}(\gamma)$ has been modified by [3] in order to add a fluid-like regularization $G_{\rm fluid}(\sigma_{f})$ of standard deviation $\sigma_{f}$. Moreover, to ensure the mapping $\gamma$ to be diffeomorphic, an alternative update of $c$ with an exponential field has been proposed in [27]: $c=s\circ\exp(u)$. The additive and diffeomorphic demons algorithms are summarized in Algotirhm 1.
@@ 9       Diffusion-like regularization: $\gamma^{[k]}=G_{\rm diff}(\sigma_{d})\star c^{[k]}$ ;
10Output: Displacement field $\gamma^{[k]}$ minimizing $E(\gamma^{[k]})$ over $k$.
Algorithm 1 Vercauteren’s demons algorithm
@@ 6       $\gamma^{[k]}$leftarrow Upsample $2^{k}\gamma^{*}$ by $2^{k}$
Algorithm 2 Multiresolution demons algorithm
@@ 4 Numerical experiments
In this section, we compare the Thirion’s demons algorithm (MATLAB built-in function imregdemons) and the additive and diffeomorphic Vercauteren’s demons algorithms (MATLAB toolbox available at https://www.mathworks.com/matlabcentral/fileexchange/39194-diffeomorphic-log-demons-image-registration) for the computation of estimates $\widetilde{\gamma}_{n}$ of the mappings $\gamma_{n}$. We consider the toy image and the sets $\Omega_{n}$ obtained by either Watershed or Voronoi partitioning presented in Figure 1.
@@ 4.1 Mapping estimation set-up
For the three demons algorithms, the diffusion-like parameter $\sigma_{d}$ and number of resolution levels $N_{\rm level}$ are selected by minimizing the quadratic risk $\|\Lambda-\Omega_{n}\circ\widetilde{\gamma}_{n}\|_{2}^{2}$ over the values $(\sigma_{d},N_{level})\in(0.3,0.31,\ldots,0.5)\times(n_{P}-1,n_{P})$, with $n_{P}$ being the highest integer such that $2^{n_{P}}$ is smaller than each dimension of the image. For the Thirions’s demons algorithm, the numbers iterations at each resolution level is set to $(2^{4},\ldots,2^{N_{level}+1})$ from the highest to the lowest level. For the Vercauteren’s demons algorithm, the maximum step length is set to $\sigma_{x}=5$, the fluid-like regularization is set to $\sigma_{f}=1$, the error threshold in the stop criterion is set to $\epsilon=10^{-3}$ and the maximum number of iterations is set to $K=500$.